### v2

# The FFT Strikes Back: An Efficient Alternative to Self-Attention

Abstract
Conventional self-attention mechanisms incur quadratic complexity, limiting their scalability on long sequences. We introduce FFTNet, an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to achieve global token mixing in $\mathcal{O}(n\log n)$ time. By transforming inputs into the frequency domain, FFTNet exploits the orthogonality and energy preservation guaranteed by Parseval’s theorem to capture long-range dependencies efficiently. A learnable spectral filter and modReLU activation dynamically emphasize salient frequency components, providing a rigorous and adaptive alternative to traditional self-attention. Experiments on the Long Range Arena and ImageNet benchmarks validate our theoretical insights and demonstrate superior performance over fixed Fourier and standard attention models.
1 Introduction
Conventional self-attention mechanisms capture global interactions through explicit pairwise computations, which results in a quadratic computational complexity that can be prohibitive for long sequences. In contrast, our work introduces an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to perform global token mixing with a mathematically elegant and scalable approach.
Our method begins by transforming the input sequence into the frequency domain, where orthogonal frequency components naturally encode long-range dependencies. This not only reduces the computational complexity to $\mathcal{O}(n\log n)$ but also preserves the energy of the original signal, as ensured by Parseval’s theorem. Such a transformation facilitates efficient global interactions without the need for exhaustive pairwise comparisons.
A central innovation of our framework is the integration of a learnable spectral filter. This adaptive component modulates the Fourier coefficients based on a global context vector, enabling the model to dynamically emphasize salient frequency bands that are critical for capturing complex patterns. Furthermore, the application of nonlinear activations to both the real and imaginary parts of the filtered signal enhances the model’s expressivity, allowing it to represent higher-order interactions that go beyond the scope of linear transformations.
In essence, our adaptive spectral filtering framework combines the computational efficiency of FFT-based transformations with adaptive, context-sensitive filtering and nonlinear processing. This synthesis offers a rigorous and expressive alternative to traditional self-attention, providing a robust solution for modeling long-range dependencies in sequence data.
2 Related Work
In this section, we review existing methods aimed at improving the efficiency of sequence models. We first discuss the complexity issues inherent in self-attention (Section 2.1), then highlight Fourier-based approaches (Section 2.2) and other approximation techniques (Section 2.3). We then examine orthogonal matrix decomposition methods (Section 2.5), and finally position our adaptive spectral filtering method within this landscape (Section 2.6).
2.1 Self-Attention Complexity
The original Transformer architecture Vaswani et al. (2017) uses pairwise dot-product attention, incurring a computational and memory cost of $\mathcal{O}(n^{2})$, where $n$ is the sequence length. As $n$ grows, this quadratic complexity quickly becomes infeasible for long sequences in tasks such as language modeling and long-context document understanding.
2.2 Fourier-Based Mixing
Fourier-based approaches leverage the Fast Fourier Transform (FFT) Cooley and Tukey (1965) to achieve more efficient global mixing of tokens. FNet Lee-Thorp et al. (2022), for example, replaces the self-attention sublayer with a fixed Fourier transform, drastically lowering computational overhead. However, the use of a static transform limits its capacity to adapt to varying inputs or highlight task-specific frequency components.
2.3 Linear, Sparse, and Low-Rank Approximations
Beyond Fourier methods, several alternative strategies aim to reduce the cost of self-attention. Performer Choromanski et al. (2021) and linear transformer variants Katharopoulos et al. (2020) approximate the softmax attention matrix to achieve linear or near-linear complexity. Meanwhile, Reformer Kitaev et al. (2020), Linformer Wang et al. (2020), and BigBird Zaheer et al. (2020) employ sparse or low-rank approximations, extending the effective context length without paying the full quadratic price. Other approaches like Synthesizer Tay et al. (2020) and MLP-Mixer Tolstikhin et al. (2021) avoid explicit token-pair interactions, replacing them with fixed or learned mixing operations.
2.4 Convolution-Based Approaches
Another line of work replaces or augments self-attention with convolutional modules to capture long-range dependencies. Methods such as Hyena Poli et al. (2023) introduce a hierarchical, convolution-based framework that can scale to long input sequences while maintaining the expressive capacity of attention. Similarly, lightweight and dynamic convolutions Wu et al. (2019) have been proposed to reduce the cost associated with attention heads, learning per-feature or per-time-step kernels to approximate global interactions. Although these methods can exhibit favorable scaling compared to vanilla self-attention, they often rely on carefully designed convolutional architectures to handle very long contexts.
2.5 Orthogonal Matrix Decomposition Methods
Orthogonal (or unitary) transformations provide a powerful avenue for stable and efficient sequence modeling. A key advantage of orthogonal decompositions is their norm-preserving property, which can mitigate issues such as vanishing or exploding gradients Wisdom et al. (2016). In the context of RNNs, unitary or orthonormal recurrent weights have been shown to preserve long-term dependencies while keeping representations stable Arjovsky et al. (2016); Lezcano-Casado and Martínez-Rubio (2019). From another perspective, the discrete Fourier transform (DFT) itself is an orthonormal transformation (up to scaling) that can mix tokens globally without explicit pairwise attention.
The FFT-based approaches discussed above can be viewed as a special class of such orthonormal transforms, where the matrix is structured by the DFT. More general orthogonal transformations—whether learned or hand-crafted—have also been proposed to reduce complexity or enhance stability in modern architectures. These include fast variants of orthonormal transforms, often parameterized in ways that ensure orthogonality is preserved throughout training Lezcano-Casado and Martínez-Rubio (2019). Within Transformers, adopting orthogonal or unitary blocks has been explored to stabilize training and capture global structure, although these methods may not always achieve the same $\mathcal{O}(n\log n)$ cost as the FFT. Nonetheless, they highlight a broad paradigm wherein structured or parameterized orthonormal decompositions serve as efficient global mixing mechanisms.
2.6 Adaptive Spectral Filtering in Context
Our work diverges from both fixed Fourier-based schemes and the various attention approximations by incorporating a learnable filter in the frequency domain. This adaptive mechanism leverages the theoretical underpinnings of FFT-based transformations—including energy preservation via Parseval’s theorem—while permitting dynamic reweighting of salient frequency bands. Thus, our method maintains an $\mathcal{O}(n\log n)$ complexity yet provides richer expressivity than fixed spectral mixing approaches. In contrast to purely approximate or sparse attention mechanisms, adaptive spectral filtering offers a direct and theoretically grounded route to capture long-range dependencies efficiently.
Overall, while approaches such as FNet, Performer, and sparse transformers demonstrate that either fixed or approximate token mixing can reduce computational overhead, our adaptive spectral filtering strategy uniquely merges the efficiency of the FFT with a learnable, input-dependent spectral filter. This provides a compelling combination of scalability and adaptability, which is crucial for complex sequence modeling tasks.
3 Adaptive Spectral Filtering Method
In this section, we present an adaptive spectral filtering framework that eliminates the need for explicit pairwise interactions in global token mixing. Instead of relying on dot-product-based self-attention, we employ the discrete Fourier transform (DFT) to capture long-range dependencies efficiently. By adaptively modulating the resulting frequency components and then applying an inverse transform, our method strikes an effective balance between expressive power and computational cost, supported by strong theoretical guarantees.
3.1 Motivation
Standard attention mechanisms compute pairwise interactions between tokens, incurring a quadratic cost in sequence length. As the number of tokens grows, this approach rapidly becomes prohibitive. In contrast, the Fourier transform decomposes a sequence into frequency components, inherently encoding global interactions in $\mathcal{O}(n\log n)$ time. To enhance representational power, we introduce a learnable filter in the frequency domain, enabling the model to emphasize salient frequency bands while retaining computational efficiency.
3.2 Method Description
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be the input sequence of length $n$ and embedding dimension $d$. Our method comprises four steps:
1. Fourier Transform.
We first apply the discrete Fourier transform along the token dimension:
$\mathbf{F}=\operatorname{FFT}(\mathbf{X})\;\in\;\mathbb{C}^{n\times d}.$
This operation represents each embedding across orthogonal frequency components, enabling global interactions without explicit pairwise comparisons.
2. Adaptive Spectral Filtering.
To selectively emphasize important frequencies, we use a learnable filter. First, compute a global context vector:
$\mathbf{c}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i},$
and pass it through a multi-layer perceptron (MLP) to obtain a modulation tensor:
$\Delta\mathbf{W}=\operatorname{MLP}(\mathbf{c})\;\in\;\mathbb{R}^{n\times d}.$
We define the final filter as
$\mathbf{W}=\mathbf{W}_{\mathrm{base}}+\Delta\mathbf{W},$
where $\mathbf{W}_{\mathrm{base}}$ is a fixed base filter (often initialized to all ones). The adaptive filtering step is then
$\tilde{\mathbf{F}}=\mathbf{F}\odot\mathbf{W},$
which reweights the Fourier coefficients element-wise according to the global context.
3. Nonlinear Activation (modReLU).
To capture higher-order relationships in the complex frequency domain, we apply the modReLU Arjovsky et al. (2016) activation, defined for a complex number $z=re^{i\theta}$ (with $r=|z|$ and $\theta=\arg(z)$) as:
$\operatorname{modReLU}(z)\;=\;\begin{cases}(r+b)\,e^{i\theta},&\text{if }r+b>0,\\[6.0pt] 0,&\text{otherwise},\end{cases}$
where $b$ is a learnable bias. This operation applies a ReLU-like threshold to the magnitude while preserving the phase, making it well suited for frequency-domain representations. The element-wise modReLU is then:
$\tilde{\mathbf{F}}=\operatorname{modReLU}\bigl{(}\tilde{\mathbf{F}}\bigr{)}.$
4. Inverse Fourier Transform.
Finally, we return to the token domain via the inverse Fourier transform, retaining the real component of the reconstructed signal:
$\mathbf{Y}=\operatorname{IFFT}\bigl{(}\tilde{\mathbf{F}}\bigr{)}\;\in\;\mathbb{R}^{n\times d}.$
This yields a globally mixed representation that incorporates adaptive filtering and nonlinear transformations in the frequency domain.
3.3 Theoretical Justification for FFT over Self-Attention
Our FFT-based adaptive spectral filtering approach offers several advantages over conventional self-attention:
  • Efficient Global Mixing: By decomposing inputs into frequency components, the FFT provides global interactions in $\mathcal{O}(n\log n)$ time—much more scalable than the $\mathcal{O}(n^{2})$ of self-attention.
  • Implicit, Adaptive Attention: The adaptive spectral filter effectively learns a frequency-domain mask informed by the global context vector. This mask reweights crucial frequency bands similarly to attention weights but avoids explicit pairwise computations.
  • Greater Expressivity via Nonlinearity: While the Fourier transform itself is linear, applying modReLU to the complex coefficients enriches representational capacity. This allows the model to capture intricate, higher-order patterns that purely linear operations may miss.
  • Energy Preservation and Stability: Parseval’s theorem ensures the input signal’s norm (energy) is preserved by the FFT. Thus, our method maintains stability by avoiding accidental loss of crucial information.
By leveraging frequency-domain global mixing, learnable spectral filtering, and the modReLU nonlinear activation, our method is both theoretically grounded and computationally efficient, serving as a robust alternative to self-attention.
3.4 Computational Complexity
The central operations in our method are the Fast Fourier Transform (FFT) and its inverse, each costing $\mathcal{O}(n\log n)$ per channel. In contrast, self-attention requires $\mathcal{O}(n^{2})$ time for pairwise computations. The adaptive filtering and activation steps introduce only a linear overhead $\mathcal{O}(n)$, keeping the overall complexity at $\mathcal{O}(n\log n)$.
3.5 Summary
In essence, adaptive spectral filtering transforms the input into the frequency domain, applies a learnable nonlinear modulation (modReLU), and then inverts the transform to obtain a globally mixed representation. This provides:
  • A favorable $\mathcal{O}(n\log n)$ complexity.
  • Robust mechanisms for modeling long-range dependencies.
  • An efficient, expressive alternative to self-attention.
The code is available here: https://github.com/jacobfa/fft.
3.6 Proof of Computational Complexity
Fourier and Inverse Fourier Transforms.
For $\mathbf{X}\in\mathbb{R}^{n\times d}$, computing the FFT and IFFT requires $\mathcal{O}(n\log n)$ operations per channel, totaling $\mathcal{O}(d\cdot n\log n)$.
Adaptive Spectral Filtering.
  • Summation over $n$ tokens for the context vector $\mathbf{c}$ takes $\mathcal{O}(n)$.
  • Computing $\Delta\mathbf{W}$ from $\mathbf{c}$ via a small MLP adds $\mathcal{O}(1)$ per channel.
  • Element-wise filtering on $\mathbf{F}$ is $\mathcal{O}(n)$ per channel.
Overall, filtering costs $\mathcal{O}(d\cdot n)$.
Nonlinear Activation.
Applying modReLU on all complex elements is $\mathcal{O}(n)$ per channel, or $\mathcal{O}(d\cdot n)$ in total.
Overall Complexity.
Since $\mathcal{O}(d\cdot n\log n)$ (for FFT/IFFT) dominates $\mathcal{O}(d\cdot n)$ (for filtering and activation), the total cost is $\mathcal{O}(d\cdot n\log n)$. For most practical settings where $d$ is not far larger than $n$, this effectively behaves as $\mathcal{O}(n\log n)$.
3.7 Proofs and Theoretical Guarantees
Below, we show key theoretical properties that justify our method as an efficient surrogate for self-attention.
3.7.1 Global Mixing via Orthogonal Decomposition
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be the input and $\mathbf{F}=\operatorname{FFT}(\mathbf{X})$. A unitary DFT matrix $\mathbf{F}_{n}\in\mathbb{C}^{n\times n}$ satisfies
$\mathbf{F}_{n}^{*}\mathbf{F}_{n}=n\mathbf{I},$
where $\mathbf{F}_{n}^{*}$ is the conjugate transpose and $\mathbf{I}$ is the identity. This orthogonality preserves inner products:
$\langle\mathbf{X}_{i},\mathbf{X}_{j}\rangle\;=\;\frac{1}{n}\langle\mathbf{F}_{i},\mathbf{F}_{j}\rangle,$
implying that each token influences every frequency component. Consequently, the transformation encodes global interactions much like self-attention, without $\mathcal{O}(n^{2})$ pairwise computations.
3.7.2 Energy Preservation via Parseval’s Theorem
Parseval’s theorem states:
$\|\mathbf{X}\|_{2}^{2}\;=\;\frac{1}{n}\,\|\mathbf{F}\|_{2}^{2}.$
After adaptive filtering and activation, the output is
$\mathbf{Y}=\operatorname{IFFT}\Bigl{(}\operatorname{modReLU}\bigl{(}\mathbf{F}\odot\mathbf{W}\bigr{)}\Bigr{)}.$
Provided the filter does not excessively amplify particular frequencies, the energy of $\mathbf{Y}$ remains close to that of $\mathbf{X}$, ensuring that key information is preserved.
3.7.3 Approximation of Self-Attention Mechanism
Self-attention can be seen as a weighted sum of tokens via
$\mathbf{A}_{ij}\;=\;\frac{\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}{\sum_{j=1}^{n}\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})},$
leading to
$\mathbf{Y}_{\text{attn}}\;=\;\mathbf{A}\,\mathbf{X}.$
Our method replaces explicit pairwise interactions with frequency-domain multiplication. By the convolution theorem, this frequency-domain multiplication is equivalent to a convolution in the token domain. Under mild conditions on $\mathbf{W}$, this global, data-dependent convolution can approximate self-attention at significantly lower computational cost.
3.7.4 Role of Nonlinear Activation in Enhancing Expressivity
While the Fourier transform is inherently linear, real-world data often exhibit nonlinear patterns. By using modReLU directly on the complex coefficients, we capture higher-order interactions that would otherwise necessitate more complex mechanisms in the token domain. The phase is preserved, and a threshold is applied to the magnitude, helping the model learn highly expressive frequency-based features.
3.8 Proof of Expressivity: FFT as an Approximation of Self-Attention
Theorem 1.
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be an input sequence, and let the self-attention mechanism be
$\mathbf{Y}_{\text{attn}}=\mathbf{A}\mathbf{X},\quad\text{where}\quad\mathbf{A}_{ij}=\frac{\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}{\sum_{j=1}^{n}\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}.$
Consider the adaptive spectral filtering operation
$\mathbf{Y}=\operatorname{IFFT}\Bigl{(}\operatorname{modReLU}\bigl{(}\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}\bigr{)}\Bigr{)},$
where $\mathbf{W}$ is a learnable frequency-domain filter and $\operatorname{modReLU}$ is a complex-valued nonlinear activation. Then, under mild regularity conditions on $\mathbf{W}$ and the activation, there exists a parameterization such that
$\mathbf{Y}\approx\mathbf{Y}_{\text{attn}},$
and the presence of the nonlinear activation extends the expressive capacity beyond that of purely linear self-attention.
Proof.
We outline the argument in several steps:
1) Unitary Transformation and Energy Preservation. The DFT matrix $\mathbf{F}_{n}$ satisfies the unitary property (up to scaling),
$\mathbf{F}_{n}^{*}\mathbf{F}_{n}=n\mathbf{I},$
implying Parseval’s theorem, which preserves the norm of $\mathbf{X}$ across the frequency transform.
2) Frequency-Domain Filtering and Convolution Equivalence. The element-wise multiplication
$\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}$
corresponds to a convolution in the token domain due to the convolution theorem. Defining $\mathbf{w}=\operatorname{IFFT}(\mathbf{W})$, we get
$\operatorname{IFFT}\Bigl{(}\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}\Bigr{)}=\mathbf{X}\ast\mathbf{w}.$
3) Approximating Self-Attention via Convolution Kernels. Self-attention can be interpreted as a learnable, global aggregation function. Convolution kernels, particularly those conditioned on input features, can approximate a wide variety of such functions. Hence, suitable choices of $\mathbf{W}$ allow $\mathbf{X}\ast\mathbf{w}$ to approximate $\mathbf{Y}_{\text{attn}}$.
4) Nonlinear Activation for Enhanced Expressivity. Because data often require nonlinear modeling, we use $\operatorname{modReLU}$ on the complex coefficients:
$z=re^{i\theta}\quad\mapsto\quad\operatorname{modReLU}(z)=\begin{cases}(r+b)e^{i\theta},&\text{if }r+b>0,\\ 0,&\text{otherwise}.\end{cases}$
This nonlinearity, when applied element-wise in the frequency domain, enriches the effective convolution kernel beyond what a purely linear approach can achieve.
5) Conclusion. By choosing $\mathbf{W}$ and the bias $b$ in $\operatorname{modReLU}$ appropriately, we can approximate the self-attention operation with an FFT-based, $\mathcal{O}(n\log n)$ method while also harnessing the additional expressive potential afforded by nonlinear activation. ∎
3.9 Comparison with Convolution-Based Methods
It is important to emphasize that simply applying a global convolutional filter does not automatically equate to attention-like behavior. Although a global convolutional kernel can, in principle, span the entire sequence, it is typically a fixed or limited set of kernels shared across positions. In contrast, self-attention computes token-dependent (query-key) interactions that vary with the input sequence in a highly adaptive manner.
In our adaptive spectral filtering approach, the learnable filter $\mathbf{W}$ depends on the global context vector $\mathbf{c}$. Once $\mathbf{c}$ is extracted, the MLP-based modulation $\Delta\mathbf{W}$ alters the frequency filter in a data-dependent way, unlike most standard convolutional methods with fixed kernels. Furthermore, the $\operatorname{modReLU}$ nonlinear activation in the frequency domain can adaptively control which frequency components are kept or suppressed, effectively yielding different convolution kernels for different inputs. This leads to a form of dynamic convolution that is more akin to the adaptability seen in attention mechanisms, albeit realized entirely in the spectral domain.
In summary, while our method can be viewed through the lens of a global convolution (due to the convolution theorem relating frequency-domain multiplication to token-domain convolution), its key novelty lies in:
  • The global context-dependent filter $\mathbf{W}$.
  • The nonlinear modulation of complex coefficients via $\operatorname{modReLU}$.
  • The ability to adapt kernel shape according to the input.
These elements collectively endow our spectral filtering strategy with adaptive, attention-like capabilities that go beyond traditional convolutional architectures.
3.10 Comparison with FNet
FNet Lee-Thorp et al. (2022) also employs the DFT to mix tokens, removing all learnable parameters for mixing. However, it lacks adaptation to specific input distributions. Our method departs from that approach in key ways:
  • Adaptive Filtering: Instead of relying on a fixed DFT, we introduce a learnable filter conditioned on a global context vector, enabling input-dependent emphasis on certain frequencies.
  • Nonlinear Activation: We incorporate a complex-domain activation (modReLU) to capture higher-order phenomena that lie beyond the scope of linear transforms.
  • Strong Theoretical Basis: Our analysis uses energy preservation (Parseval’s theorem) and the convolution theorem to formally relate frequency-domain multiplication to global token mixing.
  • Practical Scalability: Although both methods achieve $\mathcal{O}(n\log n)$ complexity, our adaptive filtering and nonlinear activation introduce minimal overhead and substantially boost expressivity.
3.10.1 Computational Efficiency
Both FFT and IFFT run in $\mathcal{O}(n\log n)$ per channel. Our adaptive filtering and modReLU steps each add only $\mathcal{O}(n)$, so the overall cost remains dominated by the FFT and IFFT. This results in an efficient, scalable method for handling long sequences, standing in contrast to $\mathcal{O}(n^{2})$ self-attention mechanisms.
In summary, adaptive spectral filtering:
  • Achieves global token mixing via the Fourier transform’s orthogonality.
  • Preserves signal energy for stable representations.
  • Approximates self-attention through a frequency-domain form of dynamic convolution.
  • Employs modReLU for higher-order interactions in the complex domain.
  • Operates in $\mathcal{O}(n\log n)$ time, offering a significant advantage over $\mathcal{O}(n^{2})$ attention-based methods.
4 Experiments
We evaluate our proposed method, FFTNet, comparing it to FNet and standard self-attention-based Transformers. We present results on the Long Range Arena (LRA) benchmark Tay et al. (2021) and the ImageNet classification task. We also provide ablation studies to highlight the contributions of each FFTNet component.
4.1 Long Range Arena (LRA) Benchmark
We evaluate on six tasks in LRA: ListOps, Text, Retrieval, Image, Pathfinder, and Path-X. Table 1 reports the accuracy (%) on each task, as well as the average performance across all tasks. Our FFTNet model achieves higher accuracy on most tasks, including a 37.65% accuracy on ListOps (compared to 36.06% for the standard Transformer and 35.33% for FNet). Overall, FFTNet slightly outperforms both baselines on average.
Model ListOps Text Retrieval Image Pathfinder Path-X Avg.
Transformer 36.06 61.54 59.67 41.51 80.38 OOM 55.83
FNet 35.33 65.11 59.61 38.67 77.80 FAIL 55.32
FFTNet (ours) 37.65 66.01 60.21 42.02 80.71 83.25 58.31
Table 1: Accuracy (%) on the Long Range Arena (LRA) benchmark. OOM indicates out-of-memory and FAIL indicates the model could not process the dataset. Our FFTNet obtains the best average accuracy.
4.2 ImageNet Classification
Next, we evaluate our FFTNetViT variants on the ImageNet classification task, comparing them to standard ViT (self-attention). Table 2 presents the FLOPs, parameter counts, and both Top-1 and Top-5 accuracy for each variant (Base, Large, Huge). We omit latency from the table for clarity and show it in a separate figure. Notably, FFTNetViT often achieves lower FLOPs than ViT for comparable model sizes, while maintaining strong accuracy. In most experiments, FNet lags behind both methods on ImageNet, so we focus on comparing our approach to the stronger self-attention baseline.
Variant FFTNetViT ViT
FLOPs Params Top-1 (%) Top-5 (%) FLOPs Params Top-1 (%) Top-5 (%)
Base 22.64 76.33 79.6 94.9 36.65 86.57 79.4 94.8
Large 79.92 267.89 82.1 96.2 127.18 304.33 81.8 96.0
Huge 166.14 539.96 83.2 96.8 261.39 632.20 82.9 96.6
Table 2: Comparison of FFTNetViT vs. Standard ViT on ImageNet, grouped by variant. FLOPs are in GFLOPs and Params in millions.
Figure 1: Latency comparison of FFTNetViT vs. Standard ViT for varying batch sizes on ImageNet. FFTNetViT scales fastear than standard self-attention.
4.3 Ablation Studies
We further investigate the contributions of individual FFTNet components by conducting ablation experiments on the ImageNet classification task (Base variant). In particular, we evaluate the following variants:
  • FFTNet (full): The complete model with spectral gating, the adaptive module, and FFT-based filtering.
  • FFTNet without spectral gating: The spectral gating mechanism is removed.
  • FFTNet without adaptive module: The adaptive module is omitted.
  • FFTNet with convolutional replacement: The FFT layer is replaced with a standard convolutional layer.
Table 3 reports the Top-1 accuracy for each variant. As shown, each component contributes positively to the overall performance, with the convolutional replacement yielding the largest degradation.
Variant Top-1 Acc (%) Observation
FFTNet (full) 79.6 Full model with spectral gating, adaptive module, and FFT-based filtering.
   – without spectral gating 78.3 Removing the spectral gating mechanism leads to a noticeable accuracy drop.
   – without adaptive module 77.8 Omitting the adaptive module further reduces performance.
   – FFT replaced with convolution 77.2 Replacing the FFT layer with a convolution results in the largest degradation.
Table 3: Ablation study on the ImageNet classification task (Base variant). Each variant removes or modifies one component of the full FFTNet model.
Overall, these experiments indicate:
  • FFTNet surpasses FNet on both LRA and ImageNet, demonstrating improved accuracy and efficiency.
  • Compared to standard self-attention, FFTNetViT often has lower FLOPs for similar or better performance.
  • Ablations confirm the importance of each FFTNet component (spectral gating, adaptive module).
5 Conclusion
We presented FFTNet, a novel approach that overcomes the inherent limitations of self-attention through adaptive spectral filtering. Our method transforms inputs into the frequency domain, leveraging Fourier theory to ensure energy preservation and efficient global mixing. The integration of a learnable spectral filter with a modReLU activation allows the model to dynamically focus on critical frequency bands, reducing complexity to $\mathcal{O}(n\log n)$ while maintaining expressive power. Extensive evaluations on LRA and ImageNet confirm that FFTNet not only achieves competitive accuracy but also significantly improves computational efficiency compared to both fixed Fourier approaches and standard self-attention. These results underscore the potential of merging rigorous theoretical foundations with adaptive learning strategies for scalable sequence modeling.