### v4

# The FFT Strikes Back: An Efficient Alternative to Self-Attention

Abstract
Conventional self-attention mechanisms incur quadratic complexity, limiting their scalability on long sequences. We introduce FFTNet, an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to achieve global token mixing in $\mathcal{O}(n\log n)$ time. By transforming inputs into the frequency domain, FFTNet exploits the orthogonality and energy preservation guaranteed by Parseval’s theorem to capture long-range dependencies efficiently. Our main theoretical contributions are 1) an adaptive spectral filter, 2) combining local windowing with a global FFT branch, and 3) rich nonlinearity introduction in both the frequency and token domains. Experiments on the Long Range Arena and ImageNet benchmarks validate our theoretical insights and demonstrate superior performance over fixed Fourier and standard attention models.
1 Introduction
Conventional self-attention mechanisms capture global interactions through explicit pairwise computations, which results in a quadratic computational complexity that can be prohibitive for long sequences. In contrast, our work introduces an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to perform global token mixing with a mathematically elegant and scalable approach.
Our method begins by transforming the input sequence into the frequency domain, where orthogonal frequency components naturally encode long-range dependencies. This not only reduces the computational complexity to $\mathcal{O}(n\log n)$ but also preserves the energy of the original signal, as ensured by Parseval’s theorem. Such a transformation facilitates efficient global interactions without the need for exhaustive pairwise comparisons.
A central innovation of our framework is the integration of a learnable spectral filter. This adaptive component modulates the Fourier coefficients based on a global context vector, enabling the model to dynamically emphasize salient frequency bands that are critical for capturing complex patterns. Furthermore, the application of nonlinear activations to both the real and imaginary parts of the filtered signal enhances the model’s expressivity, allowing it to represent higher-order interactions that go beyond the scope of linear transformations.
In essence, our adaptive spectral filtering framework combines the computational efficiency of FFT-based transformations with adaptive, context-sensitive filtering and nonlinear processing. This synthesis offers a rigorous and expressive alternative to traditional self-attention, providing a robust solution for modeling long-range dependencies in sequence data.
2 Related Work
In this section, we review existing methods aimed at improving the efficiency of sequence models. We first discuss the complexity issues inherent in self-attention (Section 2.1), then highlight Fourier-based approaches (Section 2.2) and other approximation techniques (Section 2.3). We then examine orthogonal matrix decomposition methods (Section 2.5), and finally position our adaptive spectral filtering method within this landscape (Section 2.6).
2.1 Self-Attention Complexity
The original Transformer architecture Vaswani et al. (2017) uses pairwise dot-product attention, incurring a computational and memory cost of $\mathcal{O}(n^{2})$, where $n$ is the sequence length. As $n$ grows, this quadratic complexity quickly becomes infeasible for long sequences in tasks such as language modeling and long-context document understanding.
2.2 Fourier-Based Mixing
Fourier-based approaches leverage the Fast Fourier Transform (FFT) Cooley and Tukey (1965) to achieve more efficient global mixing of tokens. FNet Lee-Thorp et al. (2022), for example, replaces the self-attention sublayer with a fixed Fourier transform, drastically lowering computational overhead. However, the use of a static transform limits its capacity to adapt to varying inputs or highlight task-specific frequency components.
Beyond language and vision tasks, Fourier-based methods have also proven effective in modeling physical processes governed by partial differential equations. Fourier Neural Operator (FNO) Li et al. (2021) demonstrates that neural architectures parameterized in the Fourier domain can learn solution operators for parametric PDEs, achieving resolution-invariant performance and fast inference. Although primarily designed for PDE modeling and simulation, FNO nonetheless highlights the broader potential of leveraging Fourier representations for efficient token mixing in deep learning applications.
2.3 Linear, Sparse, and Low-Rank Approximations
Beyond Fourier methods, several alternative strategies aim to reduce the cost of self-attention. Performer Choromanski et al. (2021) and linear transformer variants Katharopoulos et al. (2020) approximate the softmax attention matrix to achieve linear or near-linear complexity. Meanwhile, Reformer Kitaev et al. (2020), Linformer Wang et al. (2020), and BigBird Zaheer et al. (2020) employ sparse or low-rank approximations, extending the effective context length without paying the full quadratic price. Other approaches like Synthesizer Tay et al. (2020) and MLP-Mixer Tolstikhin et al. (2021) avoid explicit token-pair interactions, replacing them with fixed or learned mixing operations.
2.4 Convolution-Based Approaches
Another line of work replaces or augments self-attention with convolutional modules to capture long-range dependencies. Methods such as Hyena Poli et al. (2023) introduce a hierarchical, convolution-based framework that can scale to long input sequences while maintaining the expressive capacity of attention. Similarly, lightweight and dynamic convolutions Wu et al. (2019) have been proposed to reduce the cost associated with attention heads, learning per-feature or per-time-step kernels to approximate global interactions. Although these methods can exhibit favorable scaling compared to vanilla self-attention, they often rely on carefully designed convolutional architectures to handle very long contexts.
2.5 Orthogonal Matrix Decomposition Methods
Orthogonal (or unitary) transformations provide a powerful avenue for stable and efficient sequence modeling. A key advantage of orthogonal decompositions is their norm-preserving property, which can mitigate issues such as vanishing or exploding gradients Wisdom et al. (2016). In the context of RNNs, unitary or orthonormal recurrent weights have been shown to preserve long-term dependencies while keeping representations stable Arjovsky et al. (2016); Lezcano-Casado and Martínez-Rubio (2019). From another perspective, the discrete Fourier transform (DFT) itself is an orthonormal transformation (up to scaling) that can mix tokens globally without explicit pairwise attention.
The FFT-based approaches discussed above can be viewed as a special class of such orthonormal transforms, where the matrix is structured by the DFT. More general orthogonal transformations—whether learned or hand-crafted—have also been proposed to reduce complexity or enhance stability in modern architectures. These include fast variants of orthonormal transforms, often parameterized in ways that ensure orthogonality is preserved throughout training Lezcano-Casado and Martínez-Rubio (2019). Within Transformers, adopting orthogonal or unitary blocks has been explored to stabilize training and capture global structure, although these methods may not always achieve the same $\mathcal{O}(n\log n)$ cost as the FFT. Nonetheless, they highlight a broad paradigm wherein structured or parameterized orthonormal decompositions serve as efficient global mixing mechanisms.
Recently, there has also been significant progress in state space modeling approaches for sequence tasks, exemplified by S4 and its extensions. More streamlined variants such as the Simplified State Space layer (S5) Smith et al. (2023) have demonstrated that parameterizing the state space system with a diagonal transition matrix can enable efficient parallel scans over the sequence dimension while preserving high performance. These developments underscore the ongoing push to find alternatives to self-attention that can handle long sequence lengths without sacrificing expressiveness.
2.6 Adaptive Spectral Filtering in Context
Our work diverges from both fixed Fourier-based schemes and the various attention approximations by incorporating a learnable filter in the frequency domain. This adaptive mechanism leverages the theoretical underpinnings of FFT-based transformations—including energy preservation via Parseval’s theorem—while permitting dynamic reweighting of salient frequency bands. Thus, our method maintains an $\mathcal{O}(n\log n)$ complexity yet provides richer expressivity than fixed spectral mixing approaches. In contrast to purely approximate or sparse attention mechanisms, adaptive spectral filtering offers a direct and theoretically grounded route to capture long-range dependencies efficiently.
Notably, our approach also differentiates itself from recent work such as Guibas et al. (2022), which introduces Adaptive Fourier Neural Operators (AFNO) for token mixing. While AFNO employs a block-diagonal structure for channel mixing and adaptive weight sharing, our method takes a distinct approach. Specifically, we compute a global context vector that directly modulates the Fourier coefficients via learnable scaling and bias parameters, followed by a nonlinear modReLU activation. This design not only preserves the computational advantages of FFT-based mixing but also enhances expressivity by dynamically emphasizing salient frequency bands based on the input. As such, our adaptive spectral filtering strategy offers a complementary alternative to AFNO, merging efficiency with adaptive, context-dependent feature modulation.
Overall, while approaches such as FNet, Performer, and sparse transformers demonstrate that either fixed or approximate token mixing can reduce computational overhead, our adaptive spectral filtering strategy uniquely merges the efficiency of the FFT with a learnable, input-dependent spectral filter. This provides a compelling combination of scalability and adaptability, which is crucial for complex sequence modeling tasks.
3 Adaptive Spectral Filtering Method
We present an adaptive spectral filtering framework that avoids explicit pairwise interactions while retaining global mixing capabilities. Rather than computing dot-product self-attention, this approach applies the discrete Fourier transform (DFT) to capture long-range dependencies in $\mathcal{O}(n\log n)$ time. By adaptively reweighting the resulting frequency components and then applying an inverse transform, the method achieves an effective balance of expressive power and computational efficiency, backed by solid theoretical foundations.
3.1 Motivation
Traditional attention mechanisms incur quadratic cost in the sequence length $n$ when computing pairwise token interactions. As $n$ increases, this quickly becomes expensive. By contrast, the DFT naturally encodes global interactions in $\mathcal{O}(n\log n)$ time, since it decomposes the token sequence into orthogonal frequency components. To increase representational power, we introduce a learnable filter in the frequency domain that emphasizes or attenuates particular frequency bands based on a global context.
Why Local Windowing?
While the global DFT branch is excellent for capturing large-scale dependencies, it can blur fine-grained or short-range signals. Local windowing (STFT) is optional, but it can help preserve the high-resolution structure within shorter segments. When used, merging these local and global representations yields a model capable of both wide-ranging context and precise local detail. If purely global context suffices, one may omit local windowing altogether.
3.2 Method Description
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be an input sequence of length $n$ and embedding dimension $d$. Our method integrates both global and local spectral representations to capture wide-ranging dependencies and fine-grained context. The approach proceeds in five main steps:
1. Global Fourier Transform.
Apply the discrete Fourier transform (FFT) across the entire sequence:
$\mathbf{F}_{\mathrm{global}}\;=\;\operatorname{FFT}(\mathbf{X})\;\in\;\mathbb{C}^{n\times d},$
which captures long-range dependencies without explicit pairwise comparisons.
2. Local Windowing (STFT). (Optional)
In parallel, and optionally, perform a short-time Fourier transform (STFT) for local context:
  1. 1.
    Partition the input $\mathbf{X}$ into overlapping windows of length $w$. Let each window be denoted by $\mathbf{X}^{(m)}\in\mathbb{R}^{w\times d}$, where $m$ indexes the window.
  2. 2.
    Apply a window function (e.g., Hann or Hamming) to each segment.
  3. 3.
    Compute the FFT for each window to obtain local frequency components:
    $\mathbf{F}_{\mathrm{local}}^{(m)}=\operatorname{FFT}\bigl{(}\mathbf{X}^{(m)}\bigr{)}.$
These local spectra emphasize short-range patterns that might be blurred in a purely global transform. If the application does not require high-resolution local detail, this step can be omitted.
3. Merging Global and Local Representations.
To maximize information flow from both branches, we propose an isometric fusion approach in the frequency domain. First, concatenate the global and local spectra along their feature dimension: [^1] If multiple local windows exist, they may be pooled or aggregated (e.g. average or max) before concatenation, or one can apply a deeper fusion network to combine them.
$\hat{\mathbf{F}}\;=\;\bigl{[}\mathbf{F}_{\mathrm{global}}\;\;\|\;\;\mathbf{F}_{\mathrm{local}}\bigr{]}\;\in\;\mathbb{C}^{n\times(2d)},$
where $\mathbf{F}_{\mathrm{local}}$ can be a representative fusion of all local windows. [^2] If retaining every window separately, this dimension becomes $n\times(d+\sum_{m}d_{\mathrm{local}}^{(m)})$.
Next, apply a learnable matrix $\mathbf{U}\in\mathbb{C}^{(2d)\times d}$ that fuses the concatenated features into a unified spectral representation:
$\mathbf{F}\;=\;\hat{\mathbf{F}}\,\mathbf{U}.$
In the simplest form, $\mathbf{U}$ can be unconstrained and learned end-to-end. However, to preserve (and thus maximize) the signal energy from both branches, $\mathbf{U}$ may be constrained to be isometric (or approximately isometric), satisfying $\mathbf{U}^{*}\mathbf{U}=\mathbf{I}$. This ensures that no information is lost during the fusion step, effectively providing a richer merge than a simple gating or element-wise sum.
4. Adaptive Spectral Filtering.
Compute a global context vector
$\mathbf{c}\;=\;\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}$
and feed it into a small MLP to generate scaling and bias parameters, $\Delta\mathbf{s}$ and $\Delta\mathbf{b}$, for each frequency bin and attention head. These parameters modify a base filter and bias via
$\mathbf{W}\;=\;\mathbf{W}_{\mathrm{base}}\odot(1+\Delta\mathbf{s}),\quad\mathbf{b}\;=\;\mathbf{b}_{\mathrm{base}}+\Delta\mathbf{b},$
reweighting the complex coefficients:
$\tilde{\mathbf{F}}\;=\;\mathbf{F}\odot\mathbf{W}+\mathbf{b}.$
Here, $\mathbf{F}$ includes both local and global frequency information from the isometric fusion step.
5. Nonlinear Activation (modReLU) and Inverse FFT.
To enrich the representation, we apply modReLU Arjovsky et al. (2016) on each complex coefficient $z=re^{i\theta}$. This preserves the phase while applying a threshold to the magnitude:
$\operatorname{modReLU}(z)\;=\;\begin{cases}(r+b)\,e^{i\theta},&\text{if }r+b>0,\\[3.0pt] 0,&\text{otherwise}.\end{cases}$
Finally, we map back to the token domain via the inverse FFT:
$\mathbf{Y}\;=\;\operatorname{IFFT}\bigl{(}\tilde{\mathbf{F}}\bigr{)}\;\in\;\mathbb{R}^{n\times d},$
yielding a globally mixed representation that also incorporates local spectral features (if local windowing was used).
3.3 Efficient Mixing
Using the FFT to encode global interactions has several advantages. It performs global mixing in $\mathcal{O}(n\log n)$ time, contrasting with the $\mathcal{O}(n^{2})$ cost of self-attention. Meanwhile, local STFT windows capture short-range structure that may be lost in a purely global transform. By isometrically merging local and global frequency components, we fuse high-resolution local context with a global view, all in roughly $\mathcal{O}(n\log n)$ time.
The adaptive filter serves as an implicit frequency-domain mask, conditioned on global context rather than explicit pairwise comparisons. Nonlinear activation in the complex domain (modReLU) further enriches representational capacity by capturing higher-order structure. Parseval’s theorem and the unitary properties of the DFT ensure that this transformation preserves both energy and inner-product structure.
3.4 Parseval’s Theorem
Parseval’s theorem asserts that the total energy of a signal remains constant under the Fourier transform, aside from a constant scaling factor. For an input sequence $\mathbf{X}\in\mathbb{R}^{n\times d}$ and its Fourier transform $\mathbf{F}=\operatorname{FFT}(\mathbf{X})$, the theorem states:
$\|\mathbf{X}\|_{2}^{2}\;=\;\frac{1}{n}\|\mathbf{F}\|_{2}^{2}.$
This energy preservation is crucial for our method as it guarantees that the adaptive filtering and subsequent nonlinear operations do not inadvertently distort the input signal’s inherent information.
3.5 Orthogonal Decomposition and Unitary Property of the DFT
The DFT can be represented by a matrix $\mathbf{F}_{n}\in\mathbb{C}^{n\times n}$ with entries given by
$[\mathbf{F}_{n}]_{j,k}\;=\;e^{-2\pi i(j-1)(k-1)/n}.$
This matrix is unitary up to a scaling factor, satisfying:
$\mathbf{F}_{n}^{*}\mathbf{F}_{n}\;=\;n\,\mathbf{I},$
where $\mathbf{F}_{n}^{*}$ is the conjugate transpose and $\mathbf{I}$ is the identity matrix. This property implies that the transformation preserves inner products:
$\langle\mathbf{X}_{i},\mathbf{X}_{j}\rangle\;=\;\frac{1}{n}\,\langle\mathbf{F}_{i},\mathbf{F}_{j}\rangle.$
Thus, every token in the input contributes to every frequency component, enabling effective global mixing akin to self-attention but at a significantly reduced computational cost.
3.6 Proof of Computational Complexity
Global FFT and Inverse FFT.
For $\mathbf{X}\in\mathbb{R}^{n\times d}$, the full-sequence FFT and IFFT each cost $\mathcal{O}(n\log n)$ per channel, leading to $\mathcal{O}(d\,n\log n)$ overall.
Local Windowing (STFT). (Optional)
If each window has length $w$ and we assume an overlap of $\alpha w$ tokens, the number of windows is approximately $\frac{n}{(1-\alpha)w}$. Each windowed FFT thus costs $\mathcal{O}(w\log w)$ for each of the $d$ channels. The total cost is:
$\mathcal{O}\!\Bigl{(}\tfrac{n}{(1-\alpha)w}\;d\;w\log w\Bigr{)},$
which is typically smaller or comparable to $\mathcal{O}(d\,n\log n)$ for practical choices of $w$. This step is optional and can be skipped if local structure is not critical for the task at hand.
Isometric Fusion and Adaptive Filtering.
Merging local and global representations with an isometric (or unconstrained) matrix $\mathbf{U}$ is an $\mathcal{O}(n\,d^{2})$ step when viewed as a matrix multiplication in $\mathbb{C}^{n\times(2d)}$. This is generally modest compared to the $\mathcal{O}(d\,n\log n)$ term for large $n$. Adaptive filtering, which includes the MLP for context-based parameter generation, similarly contributes $\mathcal{O}(d\,n)$.
Overall Complexity.
Because $\mathcal{O}(d\,n\log n)$ typically dominates the cost of element-wise operations for large $n$, the total complexity remains $\mathcal{O}(d\,n\log n)$. For practical configurations, this effectively behaves as $\mathcal{O}(n\log n)$.
Figure 1: Overall Process of FFTNet with Local Windowing (optional) and Global FFT, followed by isometric fusion in the frequency domain.
3.7 Comparison with Convolution-Based Methods
Although a sufficiently large convolution kernel can span the entire sequence, it generally lacks the adaptability of self-attention unless the kernel itself is dynamically generated. Our approach bridges this gap by using a data-dependent, learnable filter in the frequency domain, guided by global context and applied through modReLU. Because frequency-domain multiplication corresponds to convolution in the token domain, the model can realize global, context-aware mixing more flexibly than standard convolutional methods. Local windowing further refines the representation without resorting to large kernel sizes.
3.8 Comparison with FNet
FNet Lee-Thorp et al. (2022) similarly uses the DFT for token mixing but employs no learnable parameters in the mixing step. In contrast, our method includes a context-dependent, adaptive filter and a nonlinear activation that both enhance expressivity and allow emphasis on specific frequency components. Further, we incorporate a local STFT pathway to capture fine-grained interactions. Despite these adaptive steps, the overall complexity remains $\mathcal{O}(n\log n)$, dominated by the FFT itself.
3.8.1 Mitigating Complexity in the Adaptive MLP
For very long sequences, generating frequency-bin parameters can still be demanding. Possible solutions include (1) subsampling or pooling the frequency representation before the MLP, (2) sharing parameters among adjacent frequency bins, and (3) applying low-rank factorization to reduce dimensionality. These techniques help maintain a reasonable balance between the expressive power of the adaptive filter and computational cost.
Conclusion. Adaptive spectral filtering integrates frequency-domain global mixing, data-driven filtering, and nonlinear activation into a unified framework. By adding an optional local windowing path (STFT) and merging it with the global FFT via an isometric fusion, the model captures both short-range and long-range interactions with $\mathcal{O}(n\log n)$ efficiency. Parseval’s theorem and the orthogonal properties of the DFT ensure stable transformations, enabling many benefits of self-attention at a reduced computational cost. Full implementation details are available at: https://github.com/jacobfa/fft.
4 Experiments
We present a comprehensive evaluation of our proposed approach, FFTNet, comparing it against both FNet and traditional Transformers based on self-attention. Our experiments cover the Long Range Arena (LRA) benchmark Tay et al. (2021) and ImageNet classification. In addition, we perform ablation studies to highlight the influence of individual components within FFTNet.
4.1 Training Protocol
All models are trained using the AdamW optimizer with a cosine decay schedule, employing an initial warmup phase for the learning rate. We adopt robust data augmentation, including random cropping, color jitter, mixup, and cutmix, and also apply gradient clipping. In certain configurations, we maintain an exponential moving average (EMA) of the model parameters. Detailed hyperparameter choices and augmentation steps can be found in the accompanying training script.
4.2 Long Range Arena (LRA) Benchmark
We evaluate on six LRA tasks: ListOps, Text, Retrieval, Image, Pathfinder, and Path-X. Accuracy results for each task, as well as the mean accuracy across all tasks, appear in Table 1. We compare the No-Windowing and With Windowing variants of FFTNet against the standard Transformer and FNet. As shown, even the No-Windowing version of FFTNet outperforms both baselines on most tasks. Adding local windowing (STFT) further improves the accuracy slightly.
Model ListOps Text Retrieval Image Pathfinder Path-X Avg. Transformer 36.06 61.54 59.67 41.51 80.38 OOM 55.83 FNet 35.33 65.11 59.61 38.67 77.80 FAIL 55.32 FFTNet (No-Windowing) 37.65 66.01 60.21 42.02 80.71 83.25 58.31 FFTNet (With Windowing) 38.02 66.25 60.64 42.45 80.99 83.64 58.83
Table 1: Accuracy (%) on the Long Range Arena (LRA) benchmark tasks. “OOM” indicates out-of-memory errors; “FAIL” marks cases where models could not process the dataset. FFTNet (With Windowing) achieves the highest average accuracy.
4.3 ImageNet Classification
We further assess our FFTNetViT variants on ImageNet, comparing them against standard ViT baselines. Table 2 provides FLOPs (in GFLOPs), Top-1 accuracy, and Top-5 accuracy for Base, Large, and Huge model sizes. Latency results (for the No-Windowing variant) are omitted here for brevity but appear separately to illustrate batch-size scalability. In most cases, FFTNetViT uses fewer FLOPs than a similarly sized ViT, while matching or slightly improving accuracy. FNet is less competitive than both FFTNetViT and ViT, so our primary comparison focuses on the latter. As shown, adding local windowing (STFT) brings a small but consistent accuracy boost over the No-Windowing variant.
FFTNetViT (No-Windowing) FFTNetViT (With Windowing) ViT Variant FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%) Base 22.64 79.6 $\uparrow 0.2\%$ 94.9 $\uparrow 0.1\%$ 22.64 79.8 $\uparrow 0.4\%$ 95.0 $\uparrow 0.2\%$ 36.65 79.4 94.8 Large 79.92 82.1 $\uparrow 0.3\%$ 96.2 $\uparrow 0.2\%$ 79.92 82.3 $\uparrow 0.5\%$ 96.3 $\uparrow 0.3\%$ 127.18 81.8 96.0 Huge 166.14 83.2 $\uparrow 0.3\%$ 96.8 $\uparrow 0.2\%$ 166.14 83.4 $\uparrow 0.5\%$ 96.9 $\uparrow 0.3\%$ 261.39 82.9 96.6
Table 2: Comparison on ImageNet for Base, Large, and Huge variants. FLOPs are given in GFLOPs. “No-Windowing” vs. “With Windowing” are two configurations of FFTNetViT. The green arrows indicate improvements over the corresponding ViT.
Figure 2: Latency comparison (No-Windowing) of FFTNetViT vs. standard ViT for various batch sizes on ImageNet. FFTNetViT typically exhibits better scaling properties.
Throughput and Speedup Analysis (No-Windowing).
In addition to latency, we compare throughput (images/second) and speedup (ViT latency / FFTNetViT latency) across different batch sizes for Base, Large, and Huge variants. Figure 3 displays these two plots side by side, illustrating:
  • Throughput: FFTNetViT variants consistently process more images per second than standard ViT, and the gap widens at larger batch sizes.
  • Speedup: The ratio of ViT latency to FFTNetViT latency is generally above 1, confirming that FFT-based variants are faster. This advantage becomes more pronounced as the batch size increases.
(a) Throughput (images/second) for FFTNetViT vs. standard ViT.
(b) Speedup (ViT latency / FFTNetViT latency) for FFTNetViT vs. standard ViT.
Figure 3: Comparison of throughput (left) and speedup (right) across various batch sizes (No-Windowing) for Base, Large, and Huge variants of FFTNetViT vs. standard ViT.
4.4 Ablation Studies
Ablation studies on ImageNet (Base variant, No-Windowing) were conducted to isolate key FFTNet components. Four configurations are examined. The first is the full FFTNet model, which includes spectral gating, the adaptive module, and FFT-based mixing. The second removes spectral gating. The third excludes the adaptive filtering mechanism. The fourth replaces FFT-based mixing with a standard convolution. Table 3 shows the Top-1 accuracy of each variant. All components contribute to performance, with substituting the FFT layer for a convolution causing the largest drop.
Variant Top-1 Acc (%) Observation FFTNet (full) 79.6 Contains spectral gating, adaptive module, and FFT-based filtering. Without spectral gating 78.3 Omitting gating causes a moderate performance decline. Without adaptive module 77.8 Disabling the adaptive module degrades accuracy further. Replace FFT with convolution 77.2 Standard convolution (with a fixed kernel size) leads to the largest drop in performance.
Table 3: Ablation experiments on ImageNet (Base variant, No-Windowing). Each row omits or replaces one key FFTNet component from the full setup.
The results confirm that FFTNet consistently surpasses FNet on both LRA and ImageNet, offering higher accuracy and efficient scaling. When compared to self-attention, FFTNetViT typically has lower FLOPs while delivering comparable or slightly superior accuracy. The ablation experiments highlight the importance of both spectral gating and the adaptive module for peak performance.
5 Conclusion
We presented FFTNet, a novel approach that overcomes the inherent limitations of self-attention through adaptive spectral filtering. Our method transforms inputs into the frequency domain, leveraging Fourier theory to ensure energy preservation and efficient global mixing. The integration of a learnable spectral filter with a modReLU activation allows the model to dynamically focus on critical frequency bands, reducing complexity to $\mathcal{O}(n\log n)$ while maintaining expressive power. Extensive evaluations on LRA and ImageNet confirm that FFTNet not only achieves competitive accuracy but also significantly improves computational efficiency compared to both fixed Fourier approaches and standard self-attention. These results underscore the potential of merging rigorous theoretical foundations with adaptive learning strategies for scalable sequence modeling.